Can you provide me with detail instructions how to construct four separate pie graphs- = 10%, 50%, 70%, 90%. I will await your prompt reply. Thank you.
Victor James






5

Answer by Fombitz(32388)

The procedures are the same for each pie except for one.
Draw the circle.
Break it into 10 equal pie slices.
For 10%, shade 1 of the 10 since 1/10=10/100=10%.
For 70%, shade 7 of the 10 since 7/10=70/100=70%.
For 90%, shade 9 of the 10 since 9/10=90/100=90%.
The easier one is 50%.
Break the pie into two equal slices and shade 1 slice since 1/2=50/100=50%.